APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday December 13, 2007 (MJD 54448) ---=== OBSERVING TEAM ===--- Elena Malanushenko (day) Shannon Watters (observer) Stephanie Snedden (observer) ---=== OBSERVING PLAN ===--- Science! plus optical collimation and a pointing model ---=== OBSERVING SUMMARY ===--- -Optical collimation improved -Pointing model obtained; pointing is much improved -Two plates observed and declared done: 2424 - Legacy 2840 - MARVELS ---=== OBSERVING LOG ===--- Afternoon: --------- All equipment are checked and work properly. Spectro collimation set for -4C and did not move. Ventilation is started at 23:20Z. Night: ------ We opened in twilight with the imager mounted for optical collimation, and started a drift on 82N at the meridian. The out-of-focus donuts appeared topheavy, but applying a tilt correction for topheavy donuts did not make them more evenly illuminated after a new center was moved to. Next we tried a plus-x tilt correction, and that helped not only the x-direction illumination pattern, but also the y-direction. We then made a small adjustment to y tilt to find the best compromise for an even illumination pattern between edge and center CCDs, and calculated and moved to the new center and tilts. We then relaxed the primary, and rechecked the donuts. The illumination pattern was the same, so we edited in the changes to default.dat. However, on our final check, (after editing default.dat), we saw that the donuts were a little "left-heavy", and the collimation could probably be tweaked again. The seeing had deteriorated, so we will leave that for another night. The new default.dat primary mirror numbers: Piston: 0 X-tilt: -7.8 Y-tilt: 3.9 X-trans: 213.5 Y-trans: 427.0 We also saved new log nominal mirror values (saveNominalMigValuesForLog). 04:07Z - We mounted the engineering camera and moved on to a 40-point pointing model. It was partly cloudy, with a promise of getting better, so we decided to not do a full 72 point model. This would give us a little more time for science. Here are the results: Comparison of new and old - note the large offset: getdat [tcc.pterr]pterr_20071214.dat Read 40 data points * fix * fit coeff change value sigma 1 IE -143.01 2 IA -149.75 3 PZZ1 -130.79 4 AN +8.20 5 AW -10.59 6 CA +34.94 7 HWCA2 -0.20 8 HWSA2 -0.16 9 HNCA1 -0.98 10 HNSA1 -0.78 11 ACEC +2.09 12 ACES -0.96 13 TX -8.11 Sky RMS = 21.58 Popn SD = 21.58 Floating the terms most relevant to the altaz telescope: * use ie ia pzz1 an aw ca * fit coeff change value sigma 1 IE -3.487 -146.50 0.976 2 IA +4.614 -145.13 1.556 3 PZZ1 +3.520 -127.27 1.121 4 AN +1.293 +9.50 0.384 5 AW +0.193 -10.40 0.384 6 CA -23.661 +11.28 1.140 7 HWCA2 -0.20 8 HWSA2 -0.16 9 HNCA1 -0.98 10 HNSA1 -0.78 11 ACEC +2.09 12 ACES -0.96 13 TX -8.11 Sky RMS = 2.08 Popn SD = 2.25 Floating all terms: * use * fit coeff change value sigma 1 IE -1.921 -148.42 1.304 2 IA -0.179 -145.31 1.282 3 PZZ1 -4.300 -131.57 3.543 4 AN -0.326 +9.17 0.411 5 AW +0.062 -10.34 0.394 6 CA +0.122 +11.40 0.976 7 HWCA2 +0.928 +0.73 0.507 8 HWSA2 -0.399 -0.56 0.566 9 HNCA1 +1.345 +0.36 0.637 10 HNSA1 +0.066 -0.71 0.632 11 ACEC -0.135 +1.96 0.680 12 ACES +1.417 +0.45 0.697 13 TX +1.795 -6.32 1.480 Sky RMS = 1.70 Popn SD = 2.07 This model should serve us for a while. 06:05Z - We mounted cartridge 9, plate 2424 - high-priority Legacy GSOGTF!!! Combined with previous signal, this plate was finished with a total exposure time of 1800 s. 07:26Z - Cartridge 7, plate 2840 - MARVELS GSOGTF! This plate has had the 10-second exposure sequence done. Tonight we set up for the 20-second exposure sequence, and then clouds formed on top of us. We waited. 08:15Z - started 25 second exposures, since there was still thin cloud moving through. None of the fibers appeared to be saturated. We will consider this plate done. 08:30Z - Cartridge 2, plate 2853 - SEGUE As soon as we mounted this plate, clouds thickened. No signal obtained. 09:45Z - Closed enclosure, but left ventilation on. 11:00Z - Ventilation off, since humidity went over the limit. Doors closed. 11:46Z - Still humid and cloudy. We won't be able to reopen. endNight didn't show the idReport for run 7236, and the idReport itself didn't have the correct final RA. This was an engineering run, so we aren't too concerned, but if there are residual problems, like true New Mexicans, we will deal with them tomorrow. Note that makeRunLog completed without any error messages after the endDrift for that run. Other than that, endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7234 23:03Z 23:10Z 100 O 137.24 139.08 26 ignore 7235 23:48Z 00:15Z 100 O 148.57 155.38 39 ignore 7237 02:58Z 03:30Z 82 N -164.73 -156.75 65 engineering ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 7236 7 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 7236 109 4 14 16.900 0.141 -2.300 0.127 0.00736 180 57 7236 195 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 7237 20 4 19 19.500 0.148 0.800 0.190 0.00758 182 57 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 7236 34 1 57 360 10 -2.105 0.214 -1.419 0.152 0.00477 7236 34 1 57 360 14 -2.061 0.224 -1.205 0.062 0.00405 7236 34 1 57 360 12 -2.097 0.133 -1.025 0.114 0.00345 7236 34 1 57 360 6 -2.113 0.090 -0.902 0.104 0.00303 7236 34 1 57 360 10 -2.209 0.195 -0.961 0.162 0.00323 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n):y QA Procedures Done (y/n):y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 47849, plate -9999 ------- 23:08 47849 0.0 bias ----- sequence 47850, plate 2424 ------- 23:51 47850 10.0 flat 23:54 47851 2.0 arc 23:58 47852 3.0 arc 00:00 47853 3.0 arc ----- sequence 47854, plate -9999 ------- 00:34 47854 0.0 bias ----- sequence 47855, plate 2424 ------- 06:46 47855 10.0 flat 06:48 47856 2.0 arc 07:07 47857 901.0 target 07:25 47858 901.0 target ----- sequence 47859, plate 2840 ------- 07:40 47859 10.0 flat 07:43 47860 2.0 arc 08:17 47861 26.0 target 08:20 47862 26.0 target 08:24 47863 26.0 target 08:27 47864 26.0 target 08:30 47865 26.0 target 08:33 47866 10.0 flat 08:36 47867 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 06:53Z 9 2424 75.34 0.0009 41.60 0.0008 236.15 0.0389 1.000150 08:15Z 7 2840 206.51 0.0063 80.54 0.0012 29.92 0.0068 1.000100 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 06:52Z 9 2424 1.00015 -1480 35 50 75 41.6 0.1 17 203 -- 1.3 08:15Z 7 2840 1.00010 -986 401 100 206 80.5 -0.8 15 201 -- 1.2 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 20:34Z 37 7 12 212 (SSW) 137 - partly cloudy 21:06Z 36 10 14 221 (SW) 120 - ' 21:38Z 37 14 14 186 (S) 173 - ' 22:11Z 35 15 10 195 (SSW) 203 - ' 22:41Z 34 13 11 206 (SSW) 174 - ' 23:14Z 32 14 15 199 (SSW) 168 - ' 23:45Z 32 13 15 198 (SSW) 180 - ' 00:15Z 32 12 14 207 (SSW) 171 - ' 00:46Z 32 12 16 207 (SSW) 195 - ' 01:02Z 32 10 16 212 (SSW) 194 - ' 01:36Z 32 11 17 202 (SSW) 179 - ' 02:07Z 32 11 15 207 (SSW) 168 - ' 02:37Z 33 9 22 223 (SW) 192 - ' 03:08Z 32 11 17 210 (SSW) 195 - ' 03:39Z 33 12 23 205 (SSW) 330 - ' 04:09Z 33 13 20 214 (SW) 225 - ' 04:40Z 32 14 18 200 (SSW) 251 - ' 05:11Z 32 13 14 192 (SSW) 205 - ' 05:43Z 32 13 18 177 (S) 206 - ' 06:13Z 32 15 21 201 (SSW) 253 - ' 06:43Z 32 17 15 184 (S) 290 - mostly clear 07:16Z 32 16 21 207 (SSW) 297 - ' 07:46Z 30 20 13 211 (SSW) 473 - cloudy 08:17Z 30 20 15 202 (SSW) 430 - ' 08:48Z 31 20 14 206 (SSW) 427 - ' 09:19Z 31 21 13 186 (S) 515 - ' 09:50Z 30 23 12 196 (SSW) 528 - ' 10:20Z 30 23 18 205 (SSW) 517 - ' 10:53Z 30 24 15 218 (SW) 569 - ' 11:25Z 29 25 16 234 (SW) 718 - ' 11:58Z 28 26 15 245 (WSW) 745 - ' ---=== TELESCOPE STATUS ===--- 23:20Z - Ventilation started 00:10Z - Opened enclosure 10:45Z - Closed enclosure, ventilation on 11:00Z - Ventilation off Telescope stowed at 30 deg alt with cartridge 2 mounted. Counterweights at 270. Spectro autofill connected. No interlocks bypassed. 11:35Z - Spectro dewar weight: 264 lb; 23 psi 11:35Z - Imager dewar weight: 126 lb; 20 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_37 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2150 -5.2130 Axial B -0.2125 -0.2135 Axial C -0.0018 -0.0016 Trans D 11.6410 11.6400 Lateral E 9.9695 9.9568 Lateral F 11.5443 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 0.2 Alt.: 87.274103 MIGS TONIGHT NOMINAL Axial A 0.8021 0.7973 Axial B 0.6129 0.6041 Axial C 0.5477 0.5425 Trans D 0.5955 0.5999 GALILS Commanded: 1636959. 1644587. 1468197. -12250. -11850. Actual: 1637658. 1645862. 1469294. -12275. -11866. SETMIR VALUES SecDesOrient: 1257.00 60.00 -9.43 -50.00 240.90 SecOrient: 1256.19 60.00 -9.61 -50.01 240.42 After collimation, we saved new nominal mirror values: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2480 -5.2480 Axial B -0.2117 -0.2117 Axial C -0.0014 -0.0014 Trans D 11.2070 11.2070 Lateral E 10.1854 10.1854 Lateral F 11.7729 11.7729 GALILS Commanded: 5100. -2500. 2600. -2450. -2550. -2700. Actual: 5113. -2479. 2616. -2457. -2538. -2689. SETMIR VALUES PriDesOrient: 0.00 -7.80 3.90 213.50 427.00 PriOrient: 1.06 -7.85 3.95 213.10 427.36 SECONDARY: ---------- Focus: 0.00 Air Temp.: -0.8 Alt.: 29.989882 MIGS TONIGHT NOMINAL Axial A 0.8025 0.8025 Axial B 0.6122 0.6122 Axial C 0.5477 0.5477 Trans D 0.6007 0.6007 GALILS Commanded: 1627447. 1653085. 1470906. -5900. -7800. Actual: 1627496. 1653168. 1470885. -5887. -7780. SETMIR VALUES SecDesOrient: 1257.00 60.00 -4.00 -20.00 118.90 SecOrient: 1256.97 59.97 -4.00 -19.92 119.30 ---=== PROBLEMS IN DETAIL ===--- Bias: Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00047849.fit -5.8 -3.6 -2.5 2.3 7.5 8.6 10.7 sdR-b2-00047849.fit -8.2 -6.2 -4.4 1.2 7.3 9.8 12.0 sdR-r1-00047849.fit -6.6 -5.2 -3.6 2.4 7.9 10.0 12.2 sdR-r2-00047849.fit -7.2 -5.7 -4.0 1.7 8.0 10.1 12.8 ---------------------------- Collimation set for -4C and did not adjust.